# Uncovering faint lensed gravitational-wave signals and reprioritizing their follow-up analysis using galaxy lensing forecasts with detected counterparts

### v2

@@ Abstract
Like light, gravitational waves can be gravitationally lensed by massive astrophysical objects. For galaxy and galaxy-cluster lenses, one expects to see strong lensing — forecasted to become observable in the coming years — where the original wave is split into multiple copies with the same frequency evolution but different overall arrival times, phases, amplitudes, and signal strengths. Some of these images can be below the detection threshold and require targeted search methods, based on tailor-made template banks. These searches can be made more sensitive by using our knowledge of the typical distribution and morphology of lenses to predict the time delay, magnification, and image-type ordering of the lensed images. Here, we show that when a subset of the images is super-threshold, they can be used to construct a more constrained prediction of the arrival time of the remaining signals, enhancing our ability to identify lensing candidate signals. Our suggested method effectively reduces the list of triggers requiring follow-up and generally re-ranks the genuine counterpart higher in the lensing candidate list. Therefore, in the future, if one observes two or three lensed images, the information they provide can be used to identify their sub-threshold counterparts, thus allowing identification of additional lensed images. Finding such images would also strengthen our evidence for the event being lensed.
@@ 1 Introduction
Gravitational waves (GWs) were predicted by general relativity more than a hundred years ago (Einstein, 1916). They are ripples in space-time propagating through the Universe and require cataclysmic phenomena to produce detectable signals. After its third observation run, the LIGO-Virgo-KAGRA Scientific collaboration detected about a hundred mergers from compact binary coalescences, originating from binary black holes (BBHs), binary neutron stars (BNSs), and neutron star black hole (NSBH) binaries (Abbott et al., 2021b). These mergers were observed by the Advanced LIGO (Aasi et al., 2015) and Advanced Virgo (Acernese et al., 2015) detectors. These observations have also enabled scientists to study new black hole and neutron star populations (Abbott et al., 2021c), cosmological expanation (Abbott et al., 2021a), and general relativity in the strong field limit (Abbott et al., 2021d), among many other things. Additionally, plans to extend the detection capabilities are ongoing, with the upgrade of the existing detectors but also new detectors being developed, such as KAGRA (Somiya, 2012; Aso et al., 2013; Akutsu et al., 2019, 2020; Abbott et al., 2020) and LIGO-India (Iyer et al., 2011). The increase afforded by more sensitive detectors will lead to increased detection rate and has the potential to open the door to further scientific opportunities. One such opportunity is the detection of lensed GW signals.
@@
Similarly to light, a massive object along GWs’ travel path can lead to gravitational lensing (Ohanian, 1974; Deguchi & Watson, 1986; Wang et al., 1996; Nakamura, 1998; Takahashi & Nakamura, 2003). Depending on the lens-source geometry and the mass of the lens, one can distinguish different types of lensing. For more massive lenses and good alignment, geometric optics applies, and the GW is split into multiple individually resolvable copies with the same frequency evolution but magnified, time delayed, and potentially undergoing an overall phase shift (Dai & Venumadhav, 2017; Ezquiaga et al., 2021). This is called strong lensing, typically produced by galaxy or galaxy cluster lenses. For galaxies, these signal copies arrive minutes to months apart (Dai & Venumadhav, 2017; Ng et al., 2018; Li et al., 2018; Oguri, 2018; Wierda et al., 2021); for galaxy clusters, the copies might arrive up to years or more apart (Smith et al., 2018, 2017, 2019; Robertson et al., 2020; Ryczanowski et al., 2020). When the lens is less massive, one might observe so-called millilensing (Liu et al., 2023). This is still the geometric optics regime but the time delay is much shorter and the various images overlap in the detector band, leading to a single non-trivial GW. For even lighter lenses, the gravitational waves enter the wave optics regime, leading to frequency-dependent beating patterns (Takahashi & Nakamura, 2003; Cao et al., 2014; Lai et al., 2018; Christian et al., 2018; Jit Singh et al., 2018; Meena & Bagla, 2020; Pagano et al., 2020; Cheung et al., 2021; Kim et al., 2020; Wright & Hendry, 2022), sometimes referred to as microlensing. Over the last years, searches for lensing have been carried out, but no universally accepted evidence has been found so far (Hannuksela et al., 2019; McIsaac et al., 2020; Li et al., 2019; Dai et al., 2020; Abbott et al., 2021e, 2023; Janquart et al., 2023c).
@@
In this work, we narrow down our focus to GW signals strongly lensed by galaxies. Such events have been forecasted by several groups to become detectable in the coming years (Ng et al., 2018; Li et al., 2018; Oguri, 2018; Wierda et al., 2021). Strong lensing by galaxies would lead to two to four potentially detectable images (Dai & Venumadhav, 2017; Ezquiaga et al., 2021; Liu et al., 2021; Lo & Magana Hernandez, 2021; Janquart et al., 2021a). To search for such events, one attempts to identify GWs with a similar frequency evolution (Haris et al., 2018; Liu et al., 2021; Lo & Magana Hernandez, 2021; Janquart et al., 2021a, 2023b; Goyal et al., 2021). One method to search for lensed events is by considering GW events already detected using already-identified events (super-threshold events) (Hannuksela et al., 2019; Abbott et al., 2021e, 2023; Janquart et al., 2023c). However, such a search can potentially miss many interesting candidates. In particular, the (de)magnification of images and the change in observation conditions due to the time delay and Earth’s rotation means that some images can exist below the noise floor as sub-threshold candidates (Li et al., 2019; McIsaac et al., 2020; Dai et al., 2020). If such sub-threshold candidates are identified, it could lead to an increased confidence in detections and lead to an improved strong lensing science case. To target such events, the seminal approach is to make use of the super-threshold events to perform an additional search using template banks (see Messick et al. (2017); Sachdev et al. (2019); Hanna et al. (2020); Cannon et al. (2020); Adams et al. (2016); Aubin et al. (2021); Allen et al. (2012); Allen (2005); Dal Canton et al. (2014); Usman et al. (2016); Nitz et al. (2017); Davies et al. (2020) for details on template bank searches) that target only event triggers with similar waveforms to the super-threshold waveforms (Li et al., 2019; McIsaac et al., 2020; Dai et al., 2020). Such sub-threshold lensing searches have already been instrumental in identifying additional lensing candidates in the LVK data (McIsaac et al., 2020; Abbott et al., 2021e, 2023; Janquart et al., 2023c). The output of the search is a list of triggers to be followed up using usual analysis methods (see Janquart et al. (2023c) for an example).
@@
However, the number of potential candidate triggers scales quadratically with time, and can thus be computationally expensive. Furthermore, the list obtained is often composed of many candidates also coming from noise artefacts and unrelated signals. Therefore, it is important to identify methods that reduce false noise triggers and computational costs.
@@
In Goyal et al. (2023), the authors utilised strong galaxy lensing forecasts to re-rank candidate triggers for follow-up analysis. Indeed, an advantage offered by galaxy strong lensing is that we have good models for this situation and we can compute the expected distributions for the time delay and relative magnification of the events (Haris et al., 2018; Wierda et al., 2021; More & More, 2022). The idea is then that those candidates with time delays consistent with strong lensing predictions should be ranked higher than those that do not.
@@
In this work, we consider a scenario where we have observed two or three super-threshold images. In particular, if a subset (for example 3 out of 4 images) of the gravitational-wave images have been already identified using super-threshold searches, our predictive power for the time delays will increase. Thus, we simulate several different scenarios in which a subset of the images have already been detected and explore how predicting the time delay for the corresponding sub-threshold images allows one to improve the ability to identify additional lensing candidates.
@@
This work is structured as follows. In Sec. 2, we describe strong lensing of GWs, focusing on galaxy lensing. Then, we give an overview of the current lensing searches for super- and sub-threshold events in Sec. 3. Next, we describe how one can construct an adapted probability density for the time delay of the sub-threshold image based on the observed super-threshold ones in Sec. 4. We illustrate the working of the method in Sec. 5. Finally, we give our conclusions and future perspectives in Sec. 6.
@@ 2 Gravitational Wave Strong Lensing by Galaxies
When strong lensing occurs, GWs are split into multiple distinct copies with the same frequency evolution [^1] When higher-order modes are present and lensing leads to a specific extra phase shift, the frequency evolution can slightly be affected. —referred to as images (Takahashi & Nakamura, 2003). Some of the images are potentially observable. Compared to unlensed signals, the lensed images undergo an overall (de)magnification, a time delay, and an overall phase shift. The waveform for the $j^{th}$ lensed image ($h^{j}_{\mathrm{L}}$) is related to the unlensed waveform ($h_{\mathrm{U}}$) as (Dai & Venumadhav, 2017; Ezquiaga et al., 2021)
@@
where the tilde denotes the frequency domain, $\vec{\theta}$ represents the usual BBH source parameters, and $\vec{\Lambda}$ represents the lensing parameters ($\vec{\Lambda}=\{\mu_{j},t_{j},n_{j}\}$, with $\mu_{j}$ the magnification changing the amplitude of the wave, $t_{j}$ the extra time delay due to the change in geometrical path and the Shapiro delay (Shapiro, 1964), and $n_{j}$ the Morse factor leading to an overall phase shift). Fig. 1 gives a representation of a lensed quadruplet.
@@
Figure 1: Illustration of an example strong lensing scenario: A galaxy lens strongly lenses a compact binary object and produces four potentially observable images that appear in the gravitational-wave detectors as multiple copies of the same signal. These copies appear otherwise identical but differ in their arrival times, amplitudes, and phases, due to lensing time delay, magnification, and image inversion, respectively. A typical galaxy lens produces either three or five images of the binary compact object, but one of the central images is typically demagnified heavily such that it is neglected and only two or four images are observed, with the two type-I images arriving first and the two type-II images arriving later. Since the amplitude and arrival time of the signal copies are different, some of the signals may reside below the noise threshold. Significant effort has been taken to dig such signals out of the noise background using searches with template banks tailored to be more sensitive to lensing. One method to improve the sensitivity of these searches is by incorporating knowledge about the distribution and morphology of galaxy lenses to predict where the signals should appear in the data.
@@
Eq. (1) is obtained by considering the stationary points of the time delay function like Fermat’s principle (Takahashi & Nakamura, 2003). Therefore, only particular points are a solution, and each of these points gives a specific value for the Morse factor. A lensed image with a certain value for its Morse factor is referred to as having a certain type. Type I images have $n_{j}=0$, corresponding to a minimum of the Fermat potential. This is equivalent to having no phase shift at all. When $n_{j}=0.5$, the image is of type II and comes from a saddle point of the potential. Finally, when the image corresponds to a maximum of the potential, we have a type III image and $n_{j}=1$. Type I and III images are indistinguishable as one can be transformed into the other by applying a $\pi/2$ change in the GW polarization, which is undetectable. Conversely, for type II images, the lensing signature is detectable provided the higher-mode content is sufficiently large (Janquart et al., 2021b; Wang et al., 2021; Vijaykumar et al., 2022).
@@
When a GW is lensed by a galaxy, we expect to see two to four images with a specific image-type ordering, with image types ordered as I–I–II–II —where the last or last two type II images may not form for some lens-source configurations, and separated by days to months (Dai & Venumadhav, 2017; Ng et al., 2018; Li et al., 2018; Oguri, 2018; Ezquiaga et al., 2021; Wierda et al., 2021; More & More, 2022). In some rare cases with a very shallow profile, we can also have a faint but observable central type III image (Dahle et al., 2013; Collett et al., 2017). This ordering is valid only for galaxies. Events strongly lensed by a galaxy cluster can have more complicated lensing morpologies (Smith et al., 2018, 2017, 2019; Robertson et al., 2020; Ryczanowski et al., 2020). We do not consider the latter in this work. Thus, when a galaxy-lensed GW event is seen, we expect the type-I images to arrive first, followed by type-II images.
@@
However, because of the change in observation condition from one image to the other —possible offline detectors from one image to the other or different sensitivity to that region of the sky because of Earth rotation— and possible demagnification relative to other images, all images are not always observable. In such scenarios, one can use the observed lensed images to predict when the other counterparts should be detected. Indeed, if we see a single type I image, then the lensing hypothesis predicts that there should exist other images that could potentially be found (see the “hidden” third gravitational wave in the illustrative Fig. 1).
@@
Observing multiple images of a lensed event can make the process of predicting the arrival time of the next image more effective. For instance, if three images are observed, rudimentary lensing theory necessitates that a fourth image must exist, even if it was not identified. Furthermore, the arrival time of the fourth image can be predicted using the observed arrival times of the other images quite accurately. As an illustrative example, we show the time-delay prediction using the software LeR (Phurailatpam, 2023). The latter produces expected lensing time-delay distributions of binary black holes sampled from the latest PowerLaw+Peak population model based on O3 data together with galaxy lenses sampled from the LSST galaxy catalogue. Similarly, for two images, the arrival time of the remaining images can be predicted, albeit with a larger uncertainty. In this work, we investigate the use of such arrival time predictions to re-rank triggers coming from sub-threshold searches when at least two super-threshold events are observed.
@@
Unfortunately, it is generally not possible to tell precisely which lensed image in the series was missed. That is, for example, in a four-image scenario with the third image having been missed (Fig. 1), we cannot tell if the missed image was the third or the fourth image. Instead, we can infer only a subset of information using the image types of the detected images. Since the image types are ordered (I–I–II–II), if we observe for example the first three images, we can infer that the missed image is of type II, but we cannot tell if it was the third or the fourth such image.
@@
Furthermore, only the Morse factor difference of the missed image can be inferred from the observed images because it is often degenerate with the coalescence phase of the waveform (Ezquiaga et al., 2021). [^2] An exception is that the Morse factor can be measurable for type II images with a large higher-order mode content (Janquart et al., 2021b; Wang et al., 2021; Vijaykumar et al., 2022). Similarly, the magnification and time delay are fully degenerate with the luminosity distance and the coalescence time of the wave, respectively (Dai & Venumadhav, 2017; Ezquiaga et al., 2021). Therefore, it is common to work with the relative lensing parameters (relative Morse phases, magnifications, and arrival times) between two images instead of the absolute parameters (absolute Morse phases, magnifications, and arrival times) (Lo & Magana Hernandez, 2021; Janquart et al., 2021a).
@@
The relative lensing parameters between an image $j$ and another image $l$ are denoted $\vec{\Phi}_{lj}=\{\mu_{lj},t_{lj},n_{lj}\}$ with the relative magnification $\mu_{lj}=\mu_{l}/\mu_{j}$, the time delay $t_{lj}=t_{l}-t_{j}$, and the Morse factor difference $n_{lj}=n_{l}-n_{j}$. Naturally, since the individual Morse factors are discrete, their differences also are. In the end, the Morse factor difference —which is generally recovered (Janquart et al., 2021a, 2023b)— does not give direct access to the individual Morse factors. However, the gravitational-wave observation does give access to the Morse factor difference between signals, which is sufficient to improve the arrival time prediction for the next image. The key assumption is that one can predict the image arrival time ordering (type I–I–II–II). For example, a morse factor difference of $n_{lj}=0$ for two detected images implies that the images are of the same type and are either the first two or the last two signals. A Morse factor difference $n_{lj}=0.5$ between two signals implies that one is type I and the other type II and they are any combination of the first two and last two signals  [^3] Strictly speaking, we could also have one is type II and the other type III but type III images are very rare and therefore we neglect them in this work.. Table 1 gives a full list of Morse factor differences and their implications. In this work, we investigate the use of these measured relative lensing parameters to determine what characteristics are expected for any missed sub-threshold images. We detail our method in Sec. 4 after a brief description of current lensing searches in the next section.

### v3

@@ Abstract
Like light, gravitational waves can be gravitationally lensed by massive astrophysical objects. Strong gravitational lensing by galaxies and galaxy clusters is anticipated to become observable in the coming years. This phenomenon will manifest as multiple copies of the original wave, each exhibiting identical frequency evolution but distinct arrival times, amplitudes, and overall phases. Some of these images can be below the detection threshold and require targeted search methods, based on tailor-made template banks. These searches can be made more sensitive by using our knowledge of the typical distribution and morphology of lenses to predict the time delay, magnification, and image-type ordering of the lensed images. Here, we show that when a subset of the galaxy lensed images is super-threshold, they can be used to construct a more constrained prediction of the arrival time of the remaining signals, enhancing our ability to identify lensing candidate signals. Our suggested method effectively reduces the list of triggers requiring follow-up and generally re-ranks the genuine counterpart higher in the lensing candidate list. So, using information provided by the two or three super-threshold images, one can identify additional lensed images, also strengthening the evidence for the lensed signal hypothesis.
@@ 1 Introduction
Gravitational waves (GWs) were predicted by general relativity more than a hundred years ago (Einstein, 1916). They are ripples in space-time propagating through the Universe and require cataclysmic phenomena to produce detectable signals. After its third observation run, the LIGO-Virgo-KAGRA (LVK) Scientific collaboration detected about a hundred mergers from compact binary coalescences, originating from binary black holes (BBHs), binary neutron stars (BNSs), and neutron star black hole (NSBH) binaries (Abbott et al., 2021b). These mergers were observed by the Advanced LIGO (Aasi et al., 2015) and Advanced Virgo (Acernese et al., 2015) detectors. These observations have also enabled scientists to study black hole and neutron star populations (Abbott et al., 2021c), cosmological expansion (Abbott et al., 2021a), and general relativity in the strong field limit (Abbott et al., 2021d), among many other things. In addition, plans are ongoing to expand detection capabilities, with the upgrade of existing detectors and the development of new detectors, such as KAGRA (Somiya, 2012; Aso et al., 2013; Akutsu et al., 2019, 2020; Abbott et al., 2020) and LIGO-India (Iyer et al., 2011). The increase afforded by more sensitive detectors will lead to an increased detection rate and has the potential to open the door to further scientific opportunities. One such opportunity is the detection of lensed GW signals.
@@
Similarly to light, a massive object along a GW’s travel path can lead to gravitational lensing (Ohanian, 1974; Deguchi & Watson, 1986; Wang et al., 1996; Nakamura, 1998; Takahashi & Nakamura, 2003). Depending on the lens-source geometry and the mass of the lens, one can distinguish different types of lensing. For more massive lenses and good alignment, geometric optics applies, and the GW is split into multiple individually resolvable copies with the same frequency evolution, but magnified, time delayed, and potentially undergoing an overall phase shift (Dai & Venumadhav, 2017; Ezquiaga et al., 2021). This is called strong lensing, typically produced by galaxy or galaxy cluster lenses. For galaxies, these signal copies arrive minutes to months apart (Dai & Venumadhav, 2017; Ng et al., 2018; Li et al., 2018; Oguri, 2018; Wierda et al., 2021); for galaxy clusters, the copies might arrive up to years or more apart (Smith et al., 2018, 2017, 2019; Robertson et al., 2020; Ryczanowski et al., 2020). When the lens is less massive, one might observe so-called millilensing (Liu et al., 2023). This is still the geometric optics regime but the time delay is much shorter and the various images overlap in the detector band, leading to a single non-trivial GW. For even lighter lenses, the GWs enter the wave optics regime, leading to frequency-dependent beating patterns (Takahashi & Nakamura, 2003; Cao et al., 2014; Lai et al., 2018; Christian et al., 2018; Jit Singh et al., 2018; Meena & Bagla, 2020; Pagano et al., 2020; Cheung et al., 2021; Kim et al., 2020; Wright & Hendry, 2022), sometimes referred to as microlensing. Over the last years, searches for lensing have been carried out, but no significant evidence has been found so far (Hannuksela et al., 2019; McIsaac et al., 2020; Li et al., 2019; Dai et al., 2020; Abbott et al., 2021e, 2023; Janquart et al., 2023c).
@@
In this work, we focus on GWs stronlgy lensed by galaxies. Several independent works have predicted such events to become detectable in the coming years (Ng et al., 2018; Li et al., 2018; Oguri, 2018; Wierda et al., 2021; Xu et al., 2021). Strong lensing by galaxies would lead to two to four potentially detectable images (Dai & Venumadhav, 2017; Ezquiaga et al., 2021; Liu et al., 2021; Lo & Magana Hernandez, 2021; Janquart et al., 2021a). To search for such events, one attempts to identify GWs with a similar frequency evolution (Haris et al., 2018; Liu et al., 2021; Lo & Magana Hernandez, 2021; Janquart et al., 2021a, 2023b; Goyal et al., 2021). One method of searching for lensed events is by considering GW events already detected using already identified events (super-threshold events) (Hannuksela et al., 2019; Abbott et al., 2021e, 2023; Janquart et al., 2023c). However, such a search can potentially miss interesting candidates. In particular, the (de)magnification of images and the change in observation conditions due to the time delay and Earth’s rotation means that some images can exist below the noise floor as sub-threshold candidates (Li et al., 2019; McIsaac et al., 2020; Dai et al., 2020; Li et al., 2023). If such sub-threshold candidates are identified, it could lead to an increased confidence in detections and lead to an improved strong lensing science case. To target such events, the seminal approach is to make use of the super-threshold events to perform an additional search using template banks (see Messick et al. (2017); Sachdev et al. (2019); Hanna et al. (2020); Cannon et al. (2020); Adams et al. (2016); Aubin et al. (2021); Allen et al. (2012); Allen (2005); Dal Canton et al. (2014); Usman et al. (2016); Nitz et al. (2017); Davies et al. (2020) for details on template bank searches) that target only event triggers with waveforms similar to the super-threshold waveforms (Li et al., 2019; McIsaac et al., 2020; Dai et al., 2020). Such sub-threshold lensing searches have already been used to search for lensed counterparts of LVK signals (McIsaac et al., 2020; Abbott et al., 2021e, 2023; Janquart et al., 2023c). The output of the search is a list of triggers to be followed up using usual analysis methods (see Abbott et al. (2023) for an example).
@@
However, the number of potential candidate triggers scales quadratically with time and can thus be computationally expensive. Furthermore, the list obtained is often composed of many candidates also coming from noise artifacts or unrelated signals. Therefore, it is important to identify methods that reduce false noise triggers and computational costs.
@@
In Goyal et al. (2023), the authors utilised strong galaxy lensing forecasts to re-rank candidate triggers for follow-up analysis; an advantage offered by galaxy strong lensing is that we have good models for this situation and we can compute the expected distributions for the time delay and relative magnification of the events (Haris et al., 2018; Wierda et al., 2021; More & More, 2022). The idea is then that those candidates with time delays consistent with strong lensing predictions should be ranked higher than those that do not.
@@
In this work, we consider a scenario where we have observed two or three super-threshold images. In particular, if a subset (for example 3 out of 4 images) of the GW images have already been identified using super-threshold searches, our predictive power for the time delays increases. Thus, we simulate several different scenarios in which a subset of the images have already been detected and explore how predicting the time delay for the corresponding sub-threshold images can improve the ability to identify additional lensing candidates.
@@
This work is structured as follows. In Sec. 2, we describe strong lensing of GWs, focusing on galaxy lensing. Then, we give an overview of the current lensing searches for super- and sub-threshold events in Sec. 3. Next, we describe how one can construct an adapted probability density for the time delay of the sub-threshold image based on the observed super-threshold ones in Sec. 4. We illustrate the method in Sec. 5. Finally, we give our conclusions and future perspectives in Sec. 6.
@@ 2 Gravitational Wave Strong Lensing by Galaxies
When strong lensing occurs, GWs are split into multiple distinct, potentially detectable, copies with the same frequency evolution [^1] When higher-order modes are present and lensing leads to a specific extra phase shift, the frequency evolution can slightly be affected.—referred to as images (Takahashi & Nakamura, 2003). Compared to unlensed signals, the lensed images undergo an overall (de)magnification, a time delay, and an overall phase shift. The waveform for the $j^{th}$ lensed image ($h^{j}_{\mathrm{L}}$) is related to the unlensed waveform ($h_{\mathrm{U}}$) as (Dai & Venumadhav, 2017; Ezquiaga et al., 2021)
@@
where the tilde denotes the frequency domain, $\vec{\theta}$ represents the usual BBH source parameters, and $\vec{\Lambda}_{j}$ represents the lensing parameters ($\vec{\Lambda}_{j}=\{\mu_{j},t_{j},n_{j}\}$, with $\mu_{j}$ the magnification changing the amplitude of the wave, $t_{j}$ the extra time delay due to the change in geometrical path and the Shapiro delay (Shapiro, 1964), and $n_{j}$ the Morse factor leading to an overall phase shift). Fig. 1 shows a representation of a lensed quadruplet.
@@
Figure 1: Illustration of an example of a strong lensing scenario. A galaxy lens strongly lenses a compact binary object and produces four potentially observable images that appear in the GW detectors as multiple copies of the same signal. These copies appear otherwise identical but differ in their arrival times, amplitudes, and can potetially undergo an overall phase shift, due to lensing time delay, magnification, and possible overall discrete phase shifts, respectively. A typical galaxy lens produces either three or five images of the binary compact object, but one of the central images is typically heavily demagnified such that it is neglected and only two or four images are observed, with the two type-I images arriving first and the two type-II images arriving later. Since the amplitude and arrival time of the signal copies are different, some of the signals may reside below the noise threshold. Significant effort has been taken to dig such signals out of the noise background using searches with template banks tailored to be more sensitive to lensing. One method to improve the sensitivity of these searches is by incorporating knowledge about the distribution and morphology of galaxy lenses to predict where the signals should appear in the data.
@@
Eq. (1) is obtained by considering the stationary points of the time delay function like Fermat’s principle (Takahashi & Nakamura, 2003). Therefore, only particular points are a solution, and each of these points gives a specific value for the Morse factor. A lensed image can be classified into different image types according to its Morse factor. Type I images have $n_{j}=0$, corresponding to a minimum of the Fermat potential. This is equivalent to having no phase shift at all. When $n_{j}=0.5$, the image is type II and comes from a saddle point of the potential. Finally, when the image corresponds to a maximum of the potential, we have a type III image and $n_{j}=1$. Type I and III images are indistinguishable as one can be transformed into the other by applying a $\pi/2$ change in the GW polarization, which is undetectable (Ezquiaga et al., 2021). In contrast, for type II images, the lensing signature is detectable provided the higher-mode content is sufficiently large (Ezquiaga et al., 2021; Janquart et al., 2021b; Wang et al., 2021; Vijaykumar et al., 2022).
@@
When a GW is lensed by a galaxy, we expect to see two to four images with a specific image-type ordering, with image types ordered as I–I–II–II —where the last or last two type II images may not form for some lens-source configurations, and separated by days to months (Dai & Venumadhav, 2017; Ng et al., 2018; Li et al., 2018; Oguri, 2018; Ezquiaga et al., 2021; Wierda et al., 2021; More & More, 2022). In some rare cases with a very shallow profile, we can also have a faint but observable central type III image (Dahle et al., 2013; Collett et al., 2017). This ordering is valid only for galaxy lenses. Events strongly lensed by a galaxy cluster can have more complicated lensing morpologies (Smith et al., 2018, 2017, 2019; Robertson et al., 2020; Ryczanowski et al., 2020). We do not consider the latter in this work. Thus, when a galaxy-lensed GW event is seen, we expect the type-I images to arrive first, followed by type-II signals.
@@
However, because of the change in observation condition from one image to the other—possible offline detectors from one image to the other or different sensitivity to that region of the sky because of Earth rotation—and possible demagnification relative to other images, all images are not always observable. In such scenarios, the observed lensed images can be used to predict when the other counterparts should be detected. Indeed, if we assume a super-threshold image is lensed and no super-threshold counterpart is found, then the lensing hypothesis predicts that there should exist other images that could potentially be found (see the “hidden” third gravitational wave in the illustrative Fig. 1).
@@
Observing multiple images of a lensed event can make the process of predicting the arrival time of the next image more effective. For instance, if three images are observed, rudimentary lensing theory necessitates that a fourth image must exist, even if it was not identified. Furthermore, the arrival time of the fourth image can be predicted using the observed arrival times of the other images quite accurately. As an illustrative example, we show the time-delay prediction using the software LeR (Phurailatpam, 2023). The latter produces expected lensing time-delay distributions of binary black holes sampled from the latest PowerLaw+Peak population model based on O3 data together with galaxy lenses sampled from the LSST galaxy catalogue (see Sec. 4 for details). Similarly, for two images, the arrival time of the remaining images can be predicted, albeit with a larger uncertainty. In this work, we investigate the use of such arrival time predictions to re-rank triggers coming from sub-threshold searches when at least two super-threshold events are observed.
@@
Unfortunately, it is generally not possible to tell precisely which lensed image in the series was missed. That is, for example, in a four-image scenario with the third image having been missed (Fig. 1), we cannot tell if the missed image was the third or the fourth image. Instead, we can infer only a subset of information using the image types of the detected images. Since the image types are ordered (I–I–II–II), we can infer that the missed image is of type II, but we cannot tell if it was the third or the fourth such image.
@@
Furthermore, only the Morse factor difference of the missed image can be inferred from the observed images because it is often degenerate with the coalescence phase of the waveform (Ezquiaga et al., 2021). [^2] An exception is that the Morse factor can be measurable for type II images with a large higher-order mode content (Janquart et al., 2021b; Wang et al., 2021; Vijaykumar et al., 2022). Similarly, the magnification and time delay are fully degenerate with the luminosity distance and the coalescence time of the wave, respectively (Dai & Venumadhav, 2017; Ezquiaga et al., 2021). Therefore, it is common to work with the relative lensing parameters (relative Morse phases, magnifications, and arrival times) between two images instead of the absolute parameters (absolute Morse phases, magnifications, and arrival times) (e.g. Lo & Magana Hernandez (2021); Janquart et al. (2021a)).
@@
The relative lensing parameters between an image $j$ and another image $l$ are denoted $\vec{\Phi}_{lj}=\{\mu_{lj},t_{lj},n_{lj}\}$ with the relative magnification $\mu_{lj}=\mu_{l}/\mu_{j}$, the time delay $t_{lj}=t_{l}-t_{j}$, and the Morse factor difference $n_{lj}=n_{l}-n_{j}$. Naturally, since the individual Morse factors are discrete, their differences also are. In the end, the Morse factor difference—which is generally recovered (Janquart et al., 2021a, 2023b)—does not give direct access to the individual Morse factors. However, the GW observation gives access to the Morse factor difference between signals, which is sufficient to improve the arrival time prediction for the next image. The key assumption is that one can predict the image arrival time ordering (type I–I–II–II). For example, a Morse factor difference of zero for two detected images implies that the images are of the same type and are either the first two or the last two signals. A Morse factor difference $n_{lj}=0.5$ between two signals implies that one is type I and the other type II and they are any combination of the first two and last two signals  [^3] Strictly speaking, we could also have one is type II and the other type III but type III images are very rare and therefore we neglect them in this work.. Table 1 gives a full list of Morse factor differences and their implications. In this work, we investigate the use of these measured relative lensing parameters to determine what characteristics are expected for any missed sub-threshold images. We detail our method in Sec. 4 after a brief description of current lensing searches in the next section.